# Straggler-Resilient Decentralized Learning via Adaptive Asynchronous Updates

### v1

@@ 4.1 Assumptions
[…]
where $p_{i}(k)$ is the number of active neighbors that worker $i$ needs to wait at iteration $k.$ Given the Metropolis weights, ${\mathbf{P}}(k)$ are doubly stochastic, i.e., $\sum_{j=1}^{N}P_{i,j}(k)=\sum_{i=1}^{N}P_{i,j}(k)=1,\forall i\in[N],j\in[N],k.$
@@
Assumptions 1-5 are standard in the literature. Assumptions 1 and 2 are common in dynamic networks with finite number of nodes [Boyd05, tang2018communication]. Specifically, Assumption 1 guarantees the product of consensus matrices ${\mathbf{P}}(1)\cdots{\mathbf{P}}(k)$ being a doubly stochastic matrix, while Assumption 2 guarantees the product matrix is a strictly positive matrix for large $k$. We define $\boldsymbol{\Phi}_{k:s}$ as the product of consensus matrices from ${\mathbf{P}}(s)$ to ${\mathbf{P}}(k)$, i.e., $\boldsymbol{\Phi}_{k:s}\triangleq{\mathbf{P}}(s){\mathbf{P}}(s+1)\cdots{\mathbf{P}}(k)$. Under Assumption 2, [nedic2018network] proved that for any integer $\ell,n$, $\boldsymbol{\Phi}_{(\ell+n)B-1:\ell B}$ is a strictly positive matrix. In fact, every entry of it is at least $\beta^{nB}$, where $\beta$ is the smallest positive value of all consensus matrices, i.e., $\beta=\arg\min_{i,j,k}P_{i,j}(k)$ with $P_{i,j}(k)>0,\forall i,j,k$. Due to the strongly-connected graph search procedure, Assumption 2 naturally holds for DSGD-AAU and $B\leq N-1.$ Assumption 3 is widely used in convergence results of gradient methods, e.g., [nedic2009distributed, bubeck2015convex, bottou2018optimization, nedic2018network]. Assumptions 4-5 are also standard [kairouz2019advances, tang2020communication, yang2021achieving]. We use a universal bound $\varsigma$ to quantify the heterogeneity of the non-i.i.d. datasets among different workers. If $\varsigma=0$, the datasets of all workers are i.i.d. It is worth noting that we do not require a bounded gradient assumption, which is often used in distributed optimization analysis [nedic2009distributed, nedic2016stochastic].
4.2 Convergence Analysis for DSGD-AAU
@@
Theorem 1.
Choose constant learning rates $\eta\leq\min\left(\sqrt{\frac{(1-q)^{2}}{30C^{2}L^{2}N}+\frac{9N^{4}}{16}}-\frac{3N^{2}}{4},1/L\right)$, where $C:=\frac{1+\beta^{-NB}}{1-\beta^{NB}}$ and $q:=(1-\beta^{NB})^{1/NB}$. Under Assumptions 1-5, the sequence of parameter $\{{\mathbf{w}}_{j}(k),\forall j\in[N]\}$ generated by DSGD-AAU satisfies
$\frac{1}{K}\sum\limits_{k=0}^{K-1}\mathbb{E}\left\|\nabla F\left(\frac{\sum_{j=1}^{N}{\mathbf{w}}_{j}(k)}{N}\right)\right\|^{2}\leq\frac{6\left(F(\bar{{\mathbf{w}}}_{0})-F({\mathbf{w}}^{*})\right)}{\eta K}+\frac{(9L+2)\eta}{3N}\sigma^{2}_{L}+\frac{2\eta}{N}\varsigma^{2},$ (7)
@@
The right hand side of (7) consists of two parts: (i) a vanishing term ${6(F\left(\bar{{\mathbf{w}}}(0)\right)-F({\mathbf{w}}^{*}))}/{\eta K}$ that goes to zero as $K$ increases; and (ii) a constant term $\frac{(9L+2)\eta}{3N}\sigma^{2}_{L}+\frac{2\eta}{N}\varsigma^{2}$ depending on the problem instance parameters and is independent of $K$. The vanishing term’s decay rate matches that of the typical DSGD methods, and the constant term converges to zero independent of $K$ as $N$ increases.
The key to achieve the bound in (7) boils down to bound the gap between individual parameter ${\mathbf{w}}_{j}(k)$ and the average $\bar{{\mathbf{w}}}(k):={\sum_{j}{\mathbf{w}}_{j}(k)}/{N}$, which is significantly affected by asynchronous updates in (3). In addition, existing state-of-the-art asynchronous method AD-PSGD [lian2018asynchronous] assumes a predefined distribution of update frequency for all workers and the global dataset should be uniformly distributed to each worker, while neither is required by our DSGD-AAU. These make their approaches not directly applied to ours and hence necessitate different proof techniques. In particular, to bound $\mathbb{E}\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\|^{2}$, AD-PSGD assumes the spectral gap and bounded staleness of each worker’s local update, while we only need to leverage the minimum non-negative element of each consensus matrices (Assumption 2). Further differentiating our proof is that we upper bound staleness periods in DSGD-AAU by $N-1$, while existing works often assume a fixed bounded stateless period, which may not hold true in practice.
With Theorem 1, we immediately have the following convergence rate for DSGD-AAU with a proper choice of learning rate:
Corollary 1.
Let $\eta=\sqrt{{N}/{K}}$. The convergence rate of DSGD-AAU is $\mathcal{O}\left(\frac{1}{\sqrt{NK}}\right)$ when $K\geq\frac{2N}{\frac{(1-q)^{2}}{30C^{2}L^{2}N}+\frac{9N^{4}}{16}-\frac{\sqrt{3}N^{3/2}(1-q)}{CL\sqrt{40}}}$.
@@
Our consensus-based decentralized optimization method with adaptive asynchronous updates still achieves a linear speedup $\mathcal{O}\left(\frac{1}{\sqrt{NK}}\right)$ with proper learning rates as shown in Corollary 1. Although many works have achieved this convergence rate asymptotically, e.g., [lian2017can] is the first to provide a theoretical analysis of decentralized SGD with a convergence rate of $\mathcal{O}(\frac{1}{\sqrt{NK}}+{\frac{1}{K}})$, these results are only for consensus-based method with synchronous updates. To the best of our knowledge, we are the first to reveal this result for fully decentralized learning with adaptive asynchronous updates, which is non-trivial due to the large-scale distributed and heterogeneous nature of training data across workers.
Input: ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$, iterations $K$, $\{{\mathbf{w}}_{j}(0)\}_{j=1}^{N}$, empty set ${\mathcal{P}},{\mathcal{V}}$.
@@
5 Realization of DSGD-AAU
Observing that for the parameters of different workers in the decentralized optimization framework to converge, each worker needs to frequently participate in the parameter update, such that every worker’s information diffuses to all other workers in the system [nedic2009distributed, nedic2016stochastic]. This is equivalent to the fact that any two arbitrary workers exchange parameters directly or indirectly through a path as the graph dynamically changes over time. The frequency for establishing such paths is mainly affected by the computation speed of stragglers. In synchronous updates, such paths are established once per “synchronous iteration”, determined by the slowest workers (see Figure 1); while in fully asynchronous updates, such a path for any worker is established immediately in each “asynchronous iteration” when it completes its local updates (see Figure 1), suffering from stale information as aforementioned. Instead, we propose DSGD-AAU to dynamically determine the subset of neighboring workers ${\mathcal{N}}_{j}(k)$ for each worker $j\in[N]$ in each iteration $k$ to establish such paths for parameter updates.
@@ 5 Realization of DSGD-AAU
In particular, for a given communication graph ${\mathcal{G}}=\{{\mathcal{N}},{\mathcal{E}}\}$, DSGD-AAU establishes the same strongly-connected graph at each worker’s side in a fully decentralized manner such that there exists at least one path for arbitrary two nodes/workers $i,j\in[N]$. We denote the set that contains all edges of such a strongly-connected graph as ${\mathcal{P}}$ and the set for all corresponding vertices as ${\mathcal{V}}$, which is essentially equal to ${\mathcal{N}}$. Our key insight is that all edges in ${\mathcal{P}}$ have been visited by DSGD-AAU at least once, i.e., all nodes in ${\mathcal{V}}$ share information with each other. Specifically, DSGD-AAU establishes sub-graphs ${\mathcal{G}}^{\prime}:=\{{\mathcal{P}},{\mathcal{V}}\}$ to dynamically determine the number of workers involved in parameter updates at each “asynchronous iteration” in a fully decentralized manner, which requires each worker to store a local copy of ${\mathcal{P}}$ and ${\mathcal{V}}$, i.e., ${\mathcal{P}}_{j},{\mathcal{V}}_{j},\forall j\in[N]$, and then to reach a consensus on $\mathcal{P}$ and $\mathcal{V}$, i.e., ${\mathcal{P}}={\mathcal{P}}_{j},{\mathcal{V}}={\mathcal{V}}_{j},\forall j$ whenever any change occurs for arbitrary ${\mathcal{P}}_{j}$ and ${\mathcal{V}}_{j}$. When a new edge $(i,j)$ is established satisfying $(i,j)\notin{\mathcal{P}}_{j},{\mathcal{P}}_{i}$, and either $j\notin{\mathcal{V}}_{i}$ or $i\notin{\mathcal{V}}_{j}$, workers $i$ and $j$ broadcast the information $(i,j)$ to the entire network such that each worker will update her own $\mathcal{P}_{j^{\prime}},\forall j^{\prime}\in[N]$ locally. Specifically, for any worker $j^{\prime}$, whenever she receives any new edges and vertices from her neighbors in ${\mathcal{N}}_{j^{\prime}}$, worker $j^{\prime}$ broadcasts this new information to her neighbors. This information sharing process continues until no new information comes. Eventually, each worker’s local set ${\mathcal{P}}_{j},{\mathcal{V}}_{j}$ reaches a consensus, i.e., ${\mathcal{P}}={\mathcal{P}}_{j},{\mathcal{V}}={\mathcal{V}}_{j},\forall j$. This procedure also avoids deadlock issues. Once the last path $(i,j)$ is established, ${\mathcal{G}}^{\prime}:=\{{\mathcal{P}},{\mathcal{V}}\}$ is strongly connected with ${\mathcal{V}}={\mathcal{N}}$. A detailed description on how to find new edges and vertices in ${\mathcal{P}}$ and ${\mathcal{V}}$ (i.e., the Pathsearch) is provided in the appendices.
@@
In the following, we will directly use ${\mathcal{G}}^{\prime}=({\mathcal{P}},{\mathcal{V}})$ to describe DSGD-AAU for simplicity, under the assumption that each worker has reached a consensus on ${\mathcal{P}}_{j}$ and ${\mathcal{V}}_{j},\forall j\in[N]$. Our DSGD-AAU algorithm (summarized in Algorithm 2) can be described in the following. We use a virtual counter $k$ to denote the iteration counter – every new edge is established no matter on which workers will increase $k$ by $1$. Let $j_{k}$ be the first worker which completes the local gradient computation in the $k$-th iteration. At each iteration $k$, the worker $j_{k}$ performs a local SGD step (lines 2-4) followed by one step of the graph searching procedure to establish a new edge in set ${\mathcal{P}}$ (line 5). The next operation in current iteration is the gossip-average of the local parameters of workers $j_{k}$ in the set ${\mathcal{N}}_{j_{k}}(k)$ (line 6). Note that if any other worker $i_{k}$ finishes local gradient update during the Pathsearch of the fast worker $j_{k}$, it will also run Pathsearch independently until ${\mathcal{P}}$ and ${\mathcal{V}}$ get updated. Hence, worker $i_{k}$ also conduct gossip-average of the local parameters in the set ${\mathcal{N}}_{i_{k}}(k)$ (lines 7-9). When a strongly-connected graph ${\mathcal{G}}^{\prime}=({\mathcal{P}},{\mathcal{V}})$ is established, reset ${\mathcal{P}}$ and ${\mathcal{V}}$ as empty sets (line 10). An illustrative example of DSGD-AAU is presented in Figure 2.
@@
The procedures in Algorithm 2 can be executed by each worker independently, and thus enables a fully decentralized implementation. Need to mention that ID sharing requires a much lower communication overhead compared with parameters sharing among neighbors. Hence no extra overhead is introduced. The overall communication overhead for each worker is upper-bounded by ${\mathcal{O}}(2NB)$ with $B$ being the maximum number of iterations for searching a strongly-connected graph ${\mathcal{G}}^{\prime}:=\{{\mathcal{P}},{\mathcal{V}}\}$, which is smaller than $N-1$. In addition, DSGD-AAU needs a memory size no more than ${\mathcal{O}}(N)$ to store a local copy of ${\mathcal{P}}$ and ${\mathcal{V}}$ as aforementioned.
Figure 2: An illustrative example of DSGD-AAU for 4 heterogeneous workers with a fully-connected network topology. At the 1st iteration ($k=1$), worker $4$ first completes the local computation, it waits until the neighbor worker $1$ finishes the local computation. Then, workers $4$ and $1$ exchange the updates and DSGD-AAU stores nodes $1,4$ into ${\mathcal{V}}$, and edge $(1,4)$ into ${\mathcal{P}}$. This terminates the 1st iteration. At the 3rd iteration ($k=3$), worker $4$ first completes the local computation, it waits until worker $1$ completes the local computation. However, since nodes $1,4$ have been stored in ${\mathcal{V}}$ and edge $(1,4)\in{\mathcal{P}}$, worker $4$ keeps waiting until worker $2$ completes the local computation. Workers $1,2,4$ exchange updated information, and DSGD-AAU stores new worker $2$ in ${\mathcal{V}}$ and ${\mathcal{P}}=\{(1,4),(2,3),(1,2),(2,4)\}$, which ends the 3rd iteration. Since there exists a path for any arbitrary two workers in ${\mathcal{G}}^{\prime}:=\{{\mathcal{P}},{\mathcal{V}}\}$ with ${\mathcal{V}}={\mathcal{N}}$, DSGD-AAU resets ${\mathcal{V}}$ and ${\mathcal{P}}$ as empty sets and move to the next iteration.
@@
Figure 3: Training loss w.r.t. iteration for different models on non-i.i.d. CIFAR-10 with 128 workers.

### v2

@@ 4.1. Assumptions
(6) $\begin{cases}P_{i,i}(k)=1-\sum_{j\in{\mathcal{N}}_{i}(k)}P_{i,j}(k),\\ P_{i,j}(k)=\frac{1}{1+\max\{p_{i}(k),p_{j}(k)\}},\quad\text{if $j\in{\mathcal{N}}_{i}(k)$},\\ P_{i,j}(k)=0,\quad\text{otherwise},\end{cases}$
where $p_{i}(k)$ is the number of active neighbors that worker $i$ needs to wait at iteration $k.$ Given the Metropolis weights, ${\mathbf{P}}(k)$ are doubly stochastic, i.e., $\sum_{j=1}^{N}P_{i,j}(k)=\sum_{i=1}^{N}P_{i,j}(k)=1,\forall i,j,k.$
@@
Assumptions 1-5 are standard in the literature. Assumptions 1 and 2 are common in dynamic networks with finite number of nodes (Xiao et al., 2006; Tang et al., 2018a). Specifically, Assumption 1 guarantees the product of consensus matrices ${\mathbf{P}}(1)\cdots{\mathbf{P}}(k)$ being a doubly stochastic matrix, while Assumption 2 guarantees the product matrix is a strictly positive matrix for large $k$. We define $\boldsymbol{\Phi}_{k:s}$ as the product of consensus matrices from ${\mathbf{P}}(s)$ to ${\mathbf{P}}(k)$, i.e., $\boldsymbol{\Phi}_{k:s}\triangleq{\mathbf{P}}(s){\mathbf{P}}(s+1)\cdots{\mathbf{P}}(k)$. Under Assumption 2, (Nedić et al., 2018) proved that for any integer $\ell,n$, $\boldsymbol{\Phi}_{(\ell+n)B-1:\ell B}$ is a strictly positive matrix. In fact, every entry of it is at least $\beta^{nB}$, where $\beta$ is the smallest positive value of all consensus matrices, i.e., $\beta=\arg\min_{i,j,k}P_{i,j}(k)$ with $P_{i,j}(k)>0,\forall i,j,k$. Due to the strongly-connected graph search procedure, Assumption 2 naturally holds for DSGD-AAU and $B\leq N-1.$ Assumption 3 is widely used in convergence results of gradient methods (Nedic and Ozdaglar, 2009; Bubeck et al., 2015; Bottou et al., 2018; Nedić et al., 2018). Assumptions 4-5 are also standard (Kairouz et al., 2019; Tang et al., 2020; Yang et al., 2021). We use a universal bound $\varsigma$ to quantify the heterogeneity of non-i.i.d. datasets among workers. If $\varsigma=0$, the datasets of all workers are i.i.d. It is worth noting that we do not require a bounded gradient assumption, which is often used in distributed optimization analysis (Nedic and Ozdaglar, 2009; Nedić and Olshevsky, 2016).
4.2. Convergence Analysis for DSGD-AAU
@@
Theorem 4.1.
Let $\eta\leq\min\left(\sqrt{\frac{(1-q)^{2}}{30C^{2}L^{2}N}+\frac{9N^{4}}{16}}-\frac{3N^{2}}{4},1/L\right)$ be constant learning rates, where $C:=\frac{1+\beta^{-NB}}{1-\beta^{NB}}$ and $q:=(1-\beta^{NB})^{1/NB}$. Under Assumptions 1-5, the sequence of parameter $\{{\mathbf{w}}_{j}(k),\forall j\in[N]\}$ generated by DSGD-AAU satisfies
$\frac{1}{K}\sum\limits_{k=0}^{K-1}\mathbb{E}\left\|\nabla F\left(\frac{\sum_{j=1}^{N}{\mathbf{w}}_{j}(k)}{N}\right)\right\|^{2}$
(7) $\qquad\qquad\leq\frac{6\left(F(\bar{{\mathbf{w}}}_{0})-F({\mathbf{w}}^{*})\right)}{\eta K}+\frac{(9L+2)\eta}{3N}\sigma^{2}_{L}+\frac{2\eta}{N}\varsigma^{2},$
@@
The right hand side of (7) consists of two parts: (i) a vanishing term ${6(F\left(\bar{{\mathbf{w}}}(0)\right)-F({\mathbf{w}}^{*}))}/{\eta K}$ that goes to zero as $K$ increases; and (ii) a constant term $\frac{(9L+2)\eta}{3N}\sigma^{2}_{L}+\frac{2\eta}{N}\varsigma^{2}$ depending on problem instance parameters and independent of $K$. The vanishing term’s decay rate matches that of typical DSGD methods, and the constant term converges to zero independent of $K$ as $N$ increases.
The key to achieve the bound in (7) boils down to bound the gap between individual parameter ${\mathbf{w}}_{j}(k)$ and the average $\bar{{\mathbf{w}}}(k):={\sum_{j}{\mathbf{w}}_{j}(k)}/{N}$, which is significantly affected by asynchronous updates in (3). In addition, existing state-of-the-art asynchronous method AD-PSGD (Lian et al., 2018) assumes a predefined distribution of update frequency for all workers and the global dataset should be uniformly distributed to each worker, while neither is required by our DSGD-AAU. These make their approaches not directly applied to ours and hence necessitate different proof techniques. In particular, to bound $\mathbb{E}\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\|^{2}$, AD-PSGD assumes the spectral gap and bounded staleness of each worker’s local update, while we only need to leverage the minimum non-negative element of each consensus matrices (Assumption 2). Further differentiating our proof is that we upper bound staleness periods in DSGD-AAU by $N-1$, while existing works often assume a fixed bounded stateless period, which may not hold true in practice.
With Theorem 4.1, we immediately have the following convergence rate for DSGD-AAU with a proper choice of learning rate:
Corollary 4.2.
Let $\eta=\sqrt{{N}/{K}}$. The convergence rate of DSGD-AAU is $\mathcal{O}\left(\frac{1}{\sqrt{NK}}\right)$ when $K\geq 30C^{2}L^{2}N^{2}$.
@@
Our consensus-based decentralized optimization method with adaptive asynchronous updates still achieves a linear speedup $\mathcal{O}\left(\frac{1}{\sqrt{NK}}\right)$ with proper learning rates as shown in Corollary 4.2. Although many works have achieved this convergence rate asymptotically, e.g., (Lian et al., 2017) is the first to provide a theoretical analysis of decentralized SGD with a convergence rate of $\mathcal{O}(\frac{1}{\sqrt{NK}}+{\frac{1}{K}})$, these results are only for consensus-based method with synchronous updates. To our best knowledge, we are the first to reveal this result for fully decentralized learning with adaptive asynchronous updates, which is non-trivial due to the large-scale distributed and heterogeneous nature of training data across workers. An interesting observation is that the value of $C>1$ is monotonically increasing as $B$ increases. Therefore, the required iteration number $K$ will become larger if the bounded-connectivity time $B$ increases. Hence, reducing the value of $B$ can lower the required number of iterations $K$. This key observation motivates our design of DSGD-AAU in Section 5, which aims to jointly dynamically select the active workers per round and minimize the connectivity time $B$. This resolves the staleness issue in conventional asynchronous algorithms (Luo et al., 2020; Assran and Rabbat, 2020; Wang et al., 2022).
5. Realization of DSGD-AAU
Intuition. Observing that for the parameters of different workers in the decentralized optimization framework to converge, each worker needs to frequently participate in the parameter update, such that every worker’s information diffuses to all other workers in the system (Nedic and Ozdaglar, 2009; Nedić and Olshevsky, 2016). This is equivalent to the fact that any two arbitrary workers exchange parameters directly or indirectly through a path as the graph dynamically changes over time. The frequency for establishing such paths is mainly affected by the computation speed of stragglers. In synchronous updates, such paths are established once per “synchronous iteration”, determined by the slowest workers (see Figure 1); while in fully asynchronous updates, such a path for any worker is established immediately in each “asynchronous iteration” when it completes its local updates (see Figure 1), suffering from stale information as aforementioned. Instead, we propose DSGD-AAU to dynamically determine the subset of neighboring workers ${\mathcal{N}}_{j}(k)$ for each worker $j\in[N]$ in each iteration $k$ to establish such paths for parameter updates.
Input: ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$, iteration numbers $K$, initialized $\{{\mathbf{w}}_{j}(0)\}_{j=1}^{N}$, and empty set ${\mathcal{P}},{\mathcal{V}}$.
@@ 5. Realization of DSGD-AAU
Realization. In particular, for a given communication graph ${\mathcal{G}}=\{{\mathcal{N}},{\mathcal{E}}\}$, DSGD-AAU establishes the same strongly-connected graph at each worker’s side in a fully decentralized manner such that there exists at least one path for arbitrary two nodes/workers $i,j\in[N]$. We denote the set that contains all edges of such a strongly-connected graph as ${\mathcal{P}}$ and the set for all corresponding vertices as ${\mathcal{V}}$, which is essentially equal to ${\mathcal{N}}$. Our key insight is that all edges in ${\mathcal{P}}$ have been visited by DSGD-AAU at least once, i.e., all nodes in ${\mathcal{V}}$ share information with each other. Specifically, DSGD-AAU establishes sub-graphs ${\mathcal{G}}^{\prime}:=\{{\mathcal{P}},{\mathcal{V}}\}$ to dynamically determine the number of workers involved in parameter updates at each “asynchronous iteration” in a fully decentralized manner, which requires each worker to store a local copy of ${\mathcal{P}}$ and ${\mathcal{V}}$, i.e., ${\mathcal{P}}_{j},{\mathcal{V}}_{j},\forall j\in[N]$, and then to reach a consensus on $\mathcal{P}$ and $\mathcal{V}$, i.e., ${\mathcal{P}}={\mathcal{P}}_{j},{\mathcal{V}}={\mathcal{V}}_{j},\forall j$ whenever any change occurs for arbitrary ${\mathcal{P}}_{j}$ and ${\mathcal{V}}_{j}$. When a new edge $(i,j)$ is established satisfying $(i,j)\notin{\mathcal{P}}_{j},{\mathcal{P}}_{i}$, and either $j\notin{\mathcal{V}}_{i}$ or $i\notin{\mathcal{V}}_{j}$, workers $i$ and $j$ broadcast the information $(i,j)$ to the entire network such that each worker will update her own $\mathcal{P}_{j^{\prime}},\forall j^{\prime}\in[N]$ locally. Specifically, for any worker $j^{\prime}$, whenever she receives any new edges and vertices from her neighbors in ${\mathcal{N}}_{j^{\prime}}$, worker $j^{\prime}$ broadcasts this new information to her neighbors. This information sharing process continues until no new information comes. Eventually, each worker’s local set ${\mathcal{P}}_{j},{\mathcal{V}}_{j}$ reaches a consensus, i.e., ${\mathcal{P}}={\mathcal{P}}_{j},{\mathcal{V}}={\mathcal{V}}_{j},\forall j$. This procedure also avoids deadlock issues. Once the last path $(i,j)$ is established, ${\mathcal{G}}^{\prime}:=\{{\mathcal{P}},{\mathcal{V}}\}$ is strongly connected with ${\mathcal{V}}={\mathcal{N}}$. A detailed description on how to find new edges and vertices in ${\mathcal{P}}$ and ${\mathcal{V}}$ (i.e., the Pathsearch) is provided in the appendices.
@@
In the following, we will directly use ${\mathcal{G}}^{\prime}=({\mathcal{P}},{\mathcal{V}})$ to describe DSGD-AAU for simplicity, under the assumption that each worker has reached a consensus on ${\mathcal{P}}_{j}$ and ${\mathcal{V}}_{j},\forall j\in[N]$. Our DSGD-AAU algorithm (summarized in Algorithm 2) can be described in the following. We use a virtual counter $k$ to denote the iteration counter – every new edge is established no matter on which workers will increase $k$ by $1$. Let $j_{k}$ be the first worker which completes the local gradient computation in the $k$-th iteration. At each iteration $k$, the worker $j_{k}$ performs a local SGD step (lines 2-4) followed by one step of the graph searching procedure to establish a new edge in set ${\mathcal{P}}$ (line 5). The next operation in current iteration is the gossip-average of the local parameters of workers $j_{k}$ in the set ${\mathcal{N}}_{j_{k}}(k)$ (line 6). Note that if any other worker $i_{k}$ finishes local gradient update during the Pathsearch of the fast worker $j_{k}$, it will also run Pathsearch independently until ${\mathcal{P}}$ and ${\mathcal{V}}$ get updated. Hence, worker $i_{k}$ also conduct gossip-average of the local parameters in the set ${\mathcal{N}}_{i_{k}}(k)$ (lines 7-9). When a strongly-connected graph ${\mathcal{G}}^{\prime}=({\mathcal{P}},{\mathcal{V}})$ is established, reset ${\mathcal{P}}$ and ${\mathcal{V}}$ as empty sets (line 10).
Pathsearch. At each iteration $k$, we denote the worker which first completes the local parameter update as $j_{k}$. Worker $j_{k}$ keeps idle until any one of the following two situations occurs. The first case is that one of worker $j_{k}$’s neighbor $i\in{\mathcal{N}}_{j_{k}}$ completes the local parameter update. If edge $(j_{k},i)\in{\mathcal{E}}\cap(j_{k},i)\notin{\mathcal{P}}$ and worker $j_{k}\notin{\mathcal{V}}\bigcup i\notin{\mathcal{V}}$, then store edge $(j_{k},i)$ into ${\mathcal{P}}$ and nodes $\{i,j_{k}\}$ into ${\mathcal{V}}$. Once this is achieved, all workers move into the $(k+1)$-th iteration. The other case is that any other new edge $(i_{1},j_{1})$ between workers $i_{1}$ and $j_{1}$ other than $j_{k}$ is established such that edge $(i_{1},j_{1})\in{\mathcal{E}}\cap(i_{1},j_{1})\notin{\mathcal{P}}$ and node $j_{1}\notin{\mathcal{V}}\bigcup i_{1}\notin{\mathcal{V}}$. Then all workers move to the $(k+1)$-th iteration. Intuitively, for both cases we guarantee that there is at least one new worker (with new updated parameters) involved in each iteration, no matter the new worker is a neighbor or for of worker $j_{k}$. This process is repeated for each iteration, which dynamically determines the subset of the fastest workers ${\mathcal{N}}(k)$. A detailed description of Pathsearch is provided in Algorithm 3 in Appendix A.
Example. An illustrative example of DSGD-AAU for 4 heterogeneous workers with a fully-connected network topology is presented in Figure 2. At the 1st iteration ($k=1$), worker $4$ first completes the local computation, it waits until the neighbor worker $1$ finishes the local computation. Then, workers $4$ and $1$ exchange the updates and DSGD-AAU stores nodes $1,4$ into ${\mathcal{V}}$, and edge $(1,4)$ into ${\mathcal{P}}$. This terminates the 1st iteration. At the 3rd iteration ($k=3$), worker $4$ first completes the local computation, it waits until worker $1$ completes the local computation. However, since nodes $1,4$ have been stored in ${\mathcal{V}}$ and edge $(1,4)\in{\mathcal{P}}$, worker $4$ keeps waiting until worker $2$ completes the local computation. Workers $1,2,4$ exchange updated information, and DSGD-AAU stores new worker $2$ in ${\mathcal{V}}$ and ${\mathcal{P}}=\{(1,4),(2,3),(1,2),(2,4)\}$, which ends the 3rd iteration. Since there exists a path for any arbitrary two workers in ${\mathcal{G}}^{\prime}:=\{{\mathcal{P}},{\mathcal{V}}\}$ with ${\mathcal{V}}={\mathcal{N}}$, DSGD-AAU resets ${\mathcal{V}}$ and ${\mathcal{P}}$ as empty sets and move to the next iteration.
Figure 2. An illustrative example of DSGD-AAU for 4 heterogeneous workers with a fully-connected network topology.
@@
The procedures in Algorithm 2 can be executed by each worker independently, and thus enables a fully decentralized implementation. Need to mention that ID sharing requires a much lower communication overhead compared with parameters sharing among neighbors. Hence no extra overhead is introduced. The overall communication overhead for each worker is upper-bounded by ${\mathcal{O}}(2NB)$ with $B$ being the maximum number of iterations for searching a strongly-connected graph ${\mathcal{G}}^{\prime}:=\{{\mathcal{P}},{\mathcal{V}}\}$, which is smaller than $N-1$. In addition, DSGD-AAU needs a memory size no more than ${\mathcal{O}}(N)$ to store a local copy of ${\mathcal{P}}$ and ${\mathcal{V}}$ as aforementioned. One of our key contributions is the Pathsearch procedure, which is essentially used to determine when a logical iteration ends. The Pathsearch procedure ends at each iteration only when a new link is established in $\mathcal{P}$. This definition of an iteration in DSGD-AAU is different from the one in AD-PSGD where a single gradient update on whichever worker is counted as one iteration. In DSGD-AAU, an iteration involves a more complex process that includes a local SGD step, a graph searching procedure, and a gossip-average operation. Moreover, since Pathsearch guarantees at each iteration, a new worker must be involved in the parameter exchange, and all workers will participate at least once every $N-1$ iterations. This is also a key property of our proposed DSGD-AAU.
@@
Figure 3. Training loss w.r.t. iteration for different models on non-i.i.d. CIFAR-10 with 128 workers.
Figure 4. Training loss w.r.t. time for different models on non-i.i.d. CIFAR-10 with 128 workers.
6. Numerical Results
We conduct extensive experiments to validate our model and theoretical results. We implement DSGD-AAU in PyTorch (Paszke et al., 2017) on Python 3 with three NVIDIA RTX A6000 GPUs using the Network File System (NFS) and MPI backends, where files are shared among workers through NFS and communications between workers are via MPI. We relegate some experimental details, parameter settings and a comprehensive set of results to Appendix C.
Baseline. We compare DSGD-AAU with three state-of-the-art methods: AD-PSGD (Lian et al., 2018), Prague (Luo et al., 2020) and AGP (Assran and Rabbat, 2020), where workers are randomly selected for parameter averaging in AD-PSGD and Prague. Hence both suffer from stragglers in the optimization problem since they ignore slower workers.
Dataset and Model. We consider four representative models: a fully-connected neural network with 2 hidden layers (2-NN), AlexNet (Krizhevsky et al., 2012), VGG-13 (Simonyan and Zisserman, 2015) and ResNet-18 (He et al., 2016) for classification tasks using CIFAR-10 (Krizhevsky et al., 2009), MNIST (LeCun et al., 1998), and Tiny-ImageNet (Russakovsky et al., 2015) datasets, focusing on non-i.i.d. versions. For example, we use the idea in (McMahan et al., 2017) to obtain non-i.i.d. CIFAR-10. We also investigate the task of next-character prediction on the dataset of The Complete Works of William Shakespeare (McMahan et al., 2017) (Shakespeare) with an LSTM language model (Kim et al., 2016). We consider a network with $32$, $64$, $128$, $256$ workers and randomly generate a connected graph for evaluation.
In real systems, there are several reasons leading to the existing of stragglers, e.g., heterogeneous hardware, hardware failure, imbalanced data distributions among tasks and different OS effects, resource contention, etc. Since we run our experiments of different number of workers in one sever as in many other works, e.g., Li et al. (Li et al., 2020a, b) and motivated by the idea of AD-PSGD (Lian et al., 2018) and Cipar et al. (Cipar et al., 2013), we randomly select workers as stragglers in each iteration. If a worker is selected to be a straggler, then it will sleep for some time in the iteration (e.g., the sleep time could be 6x of the average one local computation time). We also investigate the impact of the sleep time and the percentage of stragglers in our ablation study, see Appendix C. Such techniques have been widely used in the literature such as AD-PSGD (Lian et al., 2018), Prague (Luo et al., 2020) and Li et al. (Li et al., 2020a, b).